An object is placed 14.3 cm to the left of a diver


Physics lens problem. please help and show work please thank you!
Answer #1

apply lens formula for diverging lens

1/f = 1/v - 1/u

- 1/12 = 1/v + 1/ 14.3

v = 6.525 cm

now the above image acts as object for lens 2

object distance of which is given by

u' = 16.8 + 6.525 = 23.325 cm

=====

b)

using lens formula for otger lens

1/26.5 = 1/ v' + 1/23.235

v' = 188.584 cm to the left of other (converging) lens

=====

comment before rate in case any doubt, will reply for sure.. goodluck